V=4/3*pi*r^3
Could I have the steps for solving for r^3? I need to get r^3 alone. And sorry, but I am a returning student, my algebra is all but forgotten, and really need the answer spelled out step-by-step.

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

V=4/3*pi*r^3
------
Multiply both sides by (3/4) to get:
pi*r^3 = (3/4)V
-------------------------
Multiply both sides by (1/pi) to get:
r^3 = (1/pi)(3/4)V
r^3 = (3/(4pi))V or (3V)/(4pi)
--------------
Cheers,
Stan H

Answer by Alan3354(69443)

V=4/3*pi*r^3
---------
Multiply by 3
3V = 4*pi*r^3
Divide by 4pi
3V/(4pi) = r^3